# AI-generated faces influence gender stereotypes and racial homogenization

### v2

@@ Abstract
Text-to-image generative AI models such as Stable Diffusion are used daily by millions worldwide. However, the extent to which these models exhibit racial and gender stereotypes is not yet fully understood. Here, we document significant biases in Stable Diffusion across six races, two genders, 32 professions, and eight attributes. Additionally, we examine the degree to which Stable Diffusion depicts individuals of the same race as being similar to one another. This analysis reveals significant racial homogenization, e.g., depicting nearly all middle eastern men as dark-skinned, bearded, and wearing a traditional headdress. We then propose novel debiasing solutions that address the above stereotypes. Finally, using a preregistered experiment, we show that being presented with inclusive AI-generated faces reduces people’s racial and gender biases, while being presented with non-inclusive ones increases such biases. This persists regardless of whether the images are labeled as AI-generated. Taken together, our findings emphasize the need to address biases and stereotypes in AI-generated content.
@@ Introduction
Artificial intelligence (AI) biases refer to the systematic and unfair preferences or prejudices embedded in AI systems, often reflecting the biases present in the data used to train these systems [1]. Such biases can perpetuate and even exacerbate societal inequalities, as the AI algorithms may inadvertently discriminate against certain groups. One glaring example of AI bias is the COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) algorithm, used by judges and parole officers in the U.S. to assess the likelihood of recidivism. This algorithm, which has been used on over a million offenders to date, has been shown to produce racially biased predictions [2, 3]. Another notable example of AI bias is found in face recognition algorithms, which have been shown to discriminate based on race and gender [4]. Amazon’s endeavor to use an algorithm to evaluate job candidates based on their resumes is yet another example of AI bias; this algorithm reinforced gender preconceptions by penalizing resumes that included phrases associated with women [5]. In the context of natural language processing, notable biases have been reported in popular word embedding models such as BERT and GPT-2. Such models often associate certain occupations or stereotypes more strongly with one gender or racial group than another [6].
@@
In this study, we focus on racial and gender stereotypes in Stable Diffusion [7], one of the most popular text-to-image generative models, used daily by millions worldwide [8]. Recent studies have demonstrated that Stable Diffusion underrepresents certain races or genders [9, 10, 11], but none of these studies proposed debiasing solutions. Moreover, none of them offered a comprehensive examination of such biases across racial groups, genders, professions, and attributes. Other studies have proposed debiasing solutions [12, 13], but these are either not automated, or are unable to generate images that adequately represent complex prompts; see the Related Work section for more details. Another form of stereotype that has been completely overlooked in the literature is when individuals of the same race are depicted as being too similar to one another, e.g., when Middle Eastern men are all depicted as being bearded and dark-skinned, or when Middle Eastern women are all depicted as wearing a traditional headcover, or “hijab”. Consequently, it remains unclear whether such stereotypes (if they exists) can be addressed by diversifying the facial features of same-race individuals. Other open questions that have not been addressed to date are whether being exposed to AI-generated faces can affect people’s racial and gender biases, and whether the AI label (i.e., the fact that the images are labeled as AI-generated rather than human-generated) plays a role in this phenomenon.
To address these questions, we start off by developing a state-of-the-art classifier to predict race and gender, allowing us to quantify biases in Stable Diffusion-generated images across six races, two genders, 32 professions, and eight attributes. We then propose a novel debiasing solution, and demonstrate its ability to outperform state-of-the-art alternatives across various benchmarks. Additionally, using a measure of image similarity, we reveal that Stable Diffusion exhibits a high degree of stereotyping, depicting individuals from certain racial backgrounds as being too similar to one another. We address this issue by proposing another novel solution, which successfully increases facial diversity across genders and racial groups. Finally, using preregistered randomized controlled trials, we show that being exposed to inclusive AI-generated faces (such as those generated by our model) can reduces people’s racial and gender biases, while exposure to non-inclusive AI-generated faces (such as those generated by Stable Diffusion) can increase these biases. Our experiment provides no evidence that the AI label plays a significant role in this phenomenon, indicating that the observed effect is entirely driven by the image’s content, irrespective of its source.
@@ Results
To examine the biases in Stable Diffusion, we start off by developing a novel classifier that predicts the race and gender of any given face image. Our classifier is composed of three stages: face detection, face embedding generation, and the classification stage; see Material and Methods for technical details. To train and validate our classifier, we utilize FairFace [14]—one of the largest publicly-available datasets of face images. FairFace specifies the race (Black, East Asian, Indian, Latinx, Middle Eastern, Southeast Asian, and White), and the gender (Female, Male) of each image. We simplified FairFace’s categorization by combining the East Asian and Southeast Asian categories into a single one (Asian); see the Materials and Methods section for more details.
We trained our classifier on FairFace’s training set, and evaluated it using FairFace’s validation set. Additionally, we benchmarked our classifier against several alternatives from the literature, namely: CLIP’s zero-shot [15], Google’s FaceNet [16] + SVM [17], FairFace’s ResNet-34 [14], EfficientNet-B7 (tuning all layers) [18], and Large Vision Transformer VIT (tuning all layers) [19]. Supplementary Note 1 summarizes the results for race and gender, respectively. These results show that, across both categories, our classifier consistently achieves state-of-the-art performance in terms of accuracy, precision, recall, and F1 score.
Examining underrepresentation in Stable Diffusion
@@
Armed with our state-of-the-art classifier, we are now ready to examine the degree to which different races and genders are represented in SDXL-generated images. To this end, we generated 10,000 images using the following racial- and gender-neutral prompt: “a photo of a person”. The distribution of the resulting images is summarized by the dashed bars in Figure 1. As can be seen, White is the most represented race (47% of images), followed by Black (33%) The remaining races are all extremely underrepresented, e.g., only 3% are Asian, and only 5% are Indian. As for gender, males are overrepresented, constituting 65% of all images. These findings mirror what has been reported by Ghosh and Caliskan [11], showing that most images generated by Stable Diffusion (v2.1) representing the prompt “a front-facing photo of a person” depict light-skinned Western men.
@@
One possible explanation behind these biases could be that SDXL is merely reflecting the biases already present in the dataset on which it was trained, namely LAION-5B [20]. To examine this possibility, we used a subset of LAION-5B [21] consisting of 88,714 images; see Material and Methods for more details. The distribution of those images is summarized by the plain bars in Figure 1. As can be seen, White is even more represented in LAION-5B than SDXL (63% vs. 47%), while all remaining races are underrepresented. As for gender, both male and female individuals are equally represented with a 50-50 split. This indicates that SDXL contains its own biases beyond just the data it was trained on, e.g., raising the percentage of Black individuals from 11% to 33%, and raising the percentage of males from 50% to 65%.
@@
One way to address the aforementioned biases is to develop a model that generates images in which all races are equally represented. To this end, we introduce a fine-tuned version of SDXL, which we call: “SDXL-Inc” (where Inc is a shorthand for Inclusion). This model was created as follows. First, we identified 32 professions, and split them into 21 that are used for fine-tuning, and 11 that are used for testing (these professions are listed in Supplementary Table 1). Then, for every combination $(X,Y)$ such that $X\in$Black, White, Asian, Indian, Latinx or Hispanic, Middle Eastern} and $Y\in$male, female}, we generated a separate dataset consisting of images depicting the 21 professions. The generation of these 12 datasets (6 races × 2 genders) was done using the following prompt: “a photo of $X$ $Y$ $Z$ looking at the camera, closeup headshot facing forward, ultra quality, sharp focus”, where $Z$ denotes one of the 21 professions. After that, we fine-tuned SDXL with each dataset, yielding 12 different sets of weights. The fine-tuning was done using LORA (Low Rank Adaptation) [22], which reduces the number of trainable parameters for the downstream task; see Materials and Methods for more details. The final model (SDXL-Inc) randomly selects one of the 12 sets of weights in response to any given prompt.
@@
Figure 1: Examining biases in LAION-5B, Stable Diffusion XL (SDXL), and our SDXL-Inc. Comparing the representation of different races and genders in a random sample of 88,714 images from the LAION-5B dataset, and in a sample of 10,000 images generated by Stable Diffusion XL using the prompt: “a photo of a person”. The same prompt is then used to generate 10,000 images by our SDXL-Inc.
@@
Let us now shift our attention to professional stereotypes in SDXL. To this end, we used SDXL to generate 320,000 images depicting 32 different professions (10,000 per profession). More specifically, for each profession, $Z$, we used the following prompt: “a photo of a $Z$, looking at the camera, closeup headshot facing forward, ultra quality, sharp focus”. Moreover, to specify the types of images that we want to avoid, we used the following negative prompt: “cartoon, anime, 3d, painting, b&w, low quality”. For each profession, we used our classifier to examine the racial- and gender- composition of the resulting images.
@@
We start off by examining racial stereotypes in each profession. Figure 2a depicts the results for the 25 professions, while Supplementary Figure 2 depicts the remaining seven professions. Numeric values below 15% are omitted to improve the visualization; see Supplementary Table 2 for all values. As can be seen, in 18 out of the 25 professions, the percentage of White faces surpasses that of all other races combined, even exceeding 80% in some cases. White also happens to be the most commonly represented race in 21 out of the 25 professions. As for the remaining four, two of them are typically low-income jobs, namely Cleaner and Security Guard, and both are mostly represented by images depicting Black individuals. Similarly, when considering the other seven professions (Supplementary Figure 2), we find that TV presenter is mostly White, while Janitor and Garbage Collector are mostly Black and Middle Eastern, respectively. These findings confirm the findings of Bianchi et al. [9], showing that prestigious, high-paying professions are often represented as White.
Having examined the racial distribution, we now shift our attention to the gender distribution. The results are depicted in Figure 2b; and the exact percentages of males and females are listed in Supplementary Table 2. As can be seen, males represent 90% of the images in 16 different professions, including Doctor, Manager, and Professor. Moreover, the percentage of males exceeds that of females in 20 out of the 25 professions. As for the remaining five, they include Nurse and Secretary—two common stereotypes of women. Similar findings were reported in the literature, showing that AI-generated images tend to associate certain jobs with men and certain other jobs with women [13, 10].
Debiasing Stable Diffusion across professions and attributes
Having examined stereotypes in terms of profession, we now turn to stereotypes in terms of the following attributes: Winner, Beautiful, Intelligent, Parent, Sibling, Terrorist, Poor, and Criminal. The results of this analysis are depicted in the upper row of Figure 3a, and the corresponding numerical values are listed in Supplementary Table 5. We find that White is extremely overrepresented in the three attributes that tend to be associated with success and attractiveness, i.e., Winner, Beautiful, and Intelligent. White is also extremely overrepresented in the two family-related attributes, i.e., Parent and Sibling. In contrast, when it comes to Terrorism, Middle Eastern are the most represented race, and none of the images depict a White individual, reinforcing existing stereotypes connecting Middle Easterners to terrorism [23]. Similarly, when it comes to crime and poverty, the majority of images depict Black individuals, reinforcing common stereotypes in this regard [24]. Similar findings were reported by Bianchi et al. [9] when analyzing Stable Diffusion (v1-4), showing that attractiveness is associated with White people, thugs are associated with Black people, and terrorists are associated with Brown men with beards.
@@
Having established that SDXL exhibits biases in terms of attributes, we now evaluate SDXL-Inc’s ability to overcome these biases. To this end, we repeated the same procedure used earlier with SDXL, but using our SDXL-Inc model instead. The results for professions are depicted in the bottom row of Figure 3a. As can be seen, the underrepresentation of different races is markedly reduced compared to SDXL, as evidenced by the substantial reduction in standard deviation ($\sigma$). Importantly, none of the eight attributes considered in this analysis were used during the fine-tuning phase, and yet SDXL-Inc was able to significantly reduce the racial biases related to them. This indicates that our solution can indeed be generalized beyond the features it was fine-tuned for.
@@
Let us now reexamine professional stereotypes to further evaluate SDXL-Inc’s debiasing capability. To this end, we selected four of the professions that were used during the fine-tuning phase (Dietitian, Manager, Pharmacist, and Pilot), as well as four professions that were not used during that phase (Accountant, Journalist, Musician, and Firefighter) to further assess the generalizability of SDXL-Inc. For each profession, Figure 3b shows the racial distribution of images using SDXL (upper row) and SDXL-Inc (lower row). As can be seen, regardless of whether the profession is Black-dominated (Musician and Firefighter) or White-dominated, our solution is able to significantly reduce racial biases. Again, the reduction in standard deviation is substantial, reaching multiple folds across all professions. Similar trends are observed when examining the remaining 24 professions considered in our analysis; see Supplementary Figure 3a.
@@
b
c
Figure 4: Results of SDXL-Div. a, For each ethnicity, the distribution of cosine similarity of images generated using SDXL (plain) and SDXL-Div (dashed). b and c, Sample images of Middle Eastern individuals generated using SDXL and SDXL-Div, respectively.
Addressing racial homogenization
@@
We aimed to develop a version of SDXL capable of generating images with greater facial diversity per race compared to the original version of SDXL. To this end, we downloaded the images hosted on the Flickr-Faces-HQ dataset, which hosts high-resolution (1024×1024) images of human faces with “considerable variation in terms of age, race, and image background” [26]. Since the dataset is unlabeled, we used our classifier to infer the race of each image; see Supplementary Table 4 for the number of images per race. The resulting labelled dataset was then used to fine-tune SDXL. Again, the fine-tuning was done using LORA [22] to reduce the number of trainable parameters for the downstream task. We call the resulting model “SDXL-Div” (where Div is a shorthand for Diversity); see Materials and Methods for technical details.
@@
To evaluate SDXL-Div’s ability to increase the facial diversity of SDXL, each of the two models (SDXL and SDXL-Div) was used to generate $\sim$10,000 images per race. Then, our racial classifier was used to obtain an embedding for each image. Finally, we computed the cosine similarity between every pair of images that have the same race and are generated by the same model, resulting in $\sim$50 million cosine similarity values per race per model. The results of this analysis are depicted in Figure 4a. As can be seen, SDXL-Div is indeed able to increase the facial diversity of SDXL, regardless of the race. The greatest difference is observed in the case of Middle Eastern individuals (dropping the mean cosine similarity from 0.61 to 0.41) and Latinx individuals (from 0.55 to 0.39). To help the reader appreciate the difference in diversity between the two models, Figures 4b and 4c provide sample images of Middle Eastern individuals generated using SDXL and SDXL-Div, respectively. As can be seen, images generated using the latter model have markedly greater facial diversity compared to those generated using the former model.
User study
@@
We ran four user studies to determine whether exposure to AI-generated faces can affect people’s racial and gender biases. We estimated that to obtain a power of 0.8 to detect a medium effect size (Cohen’s d) of 0.5 in a paired-sample comparison, a sample of 135 participants would be needed. We recruited participants on Prolific, with prescreening settings of US resident and English as native language. Additionally, we prohibited participation more than once in our experiment. These studies were preregistered at AsPredicted (https://aspredicted.org/3v3wt.pdf). All studies were approved by the Institutional Review Board at New York University Abu Dhabi under the category of Exempt or Expedited Research (HRPP-2024-53).
@@
Study 2 examines gender bias, and is similar to Study 1 apart from two modifications: (i) The professions are {accountant,math $\{\textnormal{accountant},\textnormal{math scientist},and\textnormal{tailor}\}$; (ii) after seeing all six images, participants answer the question $Q_{2}$: What percentage of $P$ in the US are men? Here, the non-inclusive images are SDXL-generated, and all depict men, while the inclusive images are generated by SDXL-Inc and half depict men while the other half depicts women. In both conditions, each of the six races is depicted in a separate image; see Supplementary Figure 9.

### v3

@@ Abstract
Text-to-image generative AI models such as Stable Diffusion are used daily by millions worldwide. However, the extent to which these models exhibit racial and gender stereotypes is not yet fully understood. Here, we document significant biases in Stable Diffusion across six races, two genders, 32 professions, and eight attributes. Additionally, we examine the degree to which Stable Diffusion depicts individuals of the same race as being similar to one another. This analysis reveals significant racial homogenization, e.g., depicting nearly all Middle Eastern men as bearded, brown-skinned, and wearing traditional attire. We then propose debiasing solutions that allow users to specify the desired distributions of race and gender when generating images while minimizing racial homogenization. Finally, using a preregistered survey experiment, we find evidence that being presented with inclusive AI-generated faces reduces people’s racial and gender biases, while being presented with non-inclusive ones increases such biases, regardless of whether the images are labeled as AI-generated. Taken together, our findings emphasize the need to address biases and stereotypes in text-to-image models.
@@ Introduction
Artificial intelligence (AI) biases refer to the systematic and unfair preferences or prejudices embedded in AI models, often reflecting the biases present in the data used to train these models [1]. Such biases can perpetuate and even exacerbate societal inequalities, as the AI algorithms may inadvertently discriminate against certain groups. One glaring example of AI bias is the COMPAS (Correctional Offender Management Profiling for Alternative Sanctions) algorithm, used by judges and parole officers in the U.S. to assess the likelihood of recidivism. This algorithm, which has been used on over a million offenders to date, has been shown to produce racially biased predictions [2, 3]. Another notable example of AI bias is found in face recognition algorithms, which have been shown to discriminate based on race and gender [4]. Amazon’s endeavor to use an algorithm to evaluate job candidates based on their resumes is yet another example of AI bias; this algorithm reinforced gender preconceptions by penalizing resumes that included phrases associated with women [5]. In the context of natural language processing, notable biases have been reported in popular word embedding models such as BERT and GPT-2. Such models often associate certain occupations or stereotypes more strongly with one gender or racial group than another [6].
@@
In this study, we focus on racial and gender stereotypes in Stable Diffusion [7], one of the most popular text-to-image generative models, used daily by millions worldwide [8]. Recent studies have demonstrated that Stable Diffusion underrepresents certain races or genders [9, 10, 11], but none of these studies proposed debiasing solutions. Moreover, none of them offered a comprehensive examination of such biases across racial groups, genders, professions, and attributes. Other studies have proposed debiasing solutions [12, 13], but these are either not automated, or are unable to generate images that adequately represent complex prompts; see the Related Work section for more details. Another form of stereotype that has been largely overlooked in the literature is when individuals of the same race are depicted as being too similar to one another. Consequently, it remains unclear whether such homogenization (if it exists) can be addressed by diversifying the facial features of same-race individuals. Other open questions that have not been addressed to date are whether being exposed to AI-generated faces can affect people’s racial and gender biases, and whether the AI label (i.e., the fact that the images are labeled as AI-generated rather than human-generated) plays a role in this phenomenon.
To address these questions, we start off by developing a classifier to predict race and gender, allowing us to quantify biases in SDXL-generated images across six races, two genders, 32 professions, and eight attributes. This broad and diverse selection of social categories was crucial for examining the presence and extent of potential biases in SDXL. The inclusion of 32 professions, in particular, enables a focused examination of occupational stereotypes, a well-documented form of societal bias [14, 15] that can be inadvertently learned and reinforced by AI models. Similarly, the eight attributes, encompassing characteristics such as ‘beautiful,’ ‘intelligent,’ ‘criminal,’ and ‘terrorist,’ were carefully chosen to explore how SDXL might be associating specific personal traits and social perceptions with different racial groups. This aspect of the analysis is particularly important given prior research that highlights the harmful impact of attributing certain qualities and judgments to individuals based solely on their race [16, 17]. Our analysis documents various biases in Stable Diffusion across professions and attributes, which echo long-standing racial and gender stereotypes. Motivated by these findings, we propose a debiasing solution called SDXL-Inc (where Inc stands for inclusive), which allows users to specify their desired distributions of race and gender when generating images. Our empirical evaluation demonstrates that our solution outperforms existing alternatives across various benchmarks.
Another critical aspect that has been largely overlooked in the literature on text-to-image models is racial homogenization. The importance of this aspect stems from its connection to “Orientalism,” which critiques the West’s propensity to depict Eastern cultures and people through simplified and homogenous imagery, often reinforcing harmful stereotypes and obscuring the complexities of individual identities [18]. For instance, if an AI model consistently generates images of Middle Eastern men with beards and traditional attire, it could reinforce a narrow and potentially misleading view of this diverse population. Such homogenization can contribute to cultural insensitivity and misrepresentation of the actual diversity within a racial or ethnic group. With this in mind, we use a measure of image similarity to quantify the degree to which SDXL-generated images of the same race resemble one another. This analysis reveals that SDXL exhibits a high degree of racial homogenization. We address this issue by proposing another solution called SDXL-Div (where Div stands for diversity), and empirically demonstrate its ability to increase the diversity of facial features when depicting individuals of a given race.
Given that millions of people use text-to-image generative AI models daily [8], it is critical to not only analyze racial and gender biases in such models, but also explore the potential impact that such biases may have on people’s perceptions. Social comparison theory [19] posits that individuals evaluate themselves by drawing comparisons with others, implying that repeated exposure to biased or limited representations can negatively impact self-esteem and belonging. Similarly, stereotype threat [20] suggests that awareness of negative stereotypes can hinder the performance of individuals from stereotyped groups. For instance, if AI models consistently portray certain professions as dominated by a particular race or gender, it could reinforce existing stereotypes and discourage others from pursuing those careers. Therefore, it is essential to investigate whether exposure to faces generated by AI can affect people’s biases, and whether the awareness of the images being AI-generated plays a role in this phenomenon. We answer this question using preregistered randomized controlled trials, the outcome of which demonstrates that being exposed to inclusive AI-generated faces reduces people’s racial and gender biases, while exposure to non-inclusive ones increases those biases. Our experiment provides no evidence that the AI label plays a significant role in this phenomenon, indicating that the observed effect is entirely driven by the image’s content, irrespective of its source.
@@ Results
To examine biases in SDXL, we developed a classifier of race and gender, composed of three stages: face detection, face embedding generation, and the classification stage; see Material and Methods for technical details. To train and validate our classifier, we utilize FairFace [21]—one of the largest publicly-available datasets of face images. FairFace specifies the race (Black, East Asian, Indian, Latinx, Middle Eastern, Southeast Asian, and White), and gender (Female, Male) of each image. These classes are widely studied in literature due to their availability from FairFace. It should be noted, however, that these classes do not capture the full spectrum of genders and races. Moreover, existing image classifiers (including our own) are meant to predict the perceived (rather than identified) gender and race. We simplified FairFace’s categorization by combining the East and Southeast Asian categories into a single one (Asian). We trained and evaluated our classifier on FairFace’s training set and validation set, respectively. Our classifier was benchmarked against several alternatives from the literature, namely: CLIP’s zero-shot [22], Google’s FaceNet [23] + SVM [24], FairFace’s ResNet-34 [21], EfficientNet-B7 (tuning all layers) [25], and Large Vision Transformer VIT (tuning all layers) [26]. Supplementary Note 1 summarizes the results, showing that our classifier consistently achieves state-of-the-art performance in terms of accuracy, precision, recall, and F1 score.
Examining biases in Stable Diffusion
@@
We examined the degree to which different races and genders appear in SDXL generated images. To this end, we generated 10,000 images using the following racial- and gender-neutral prompt: “a photo of a person”. The distribution of the resulting images is summarized by the dashed bars in Figure 1. As can be seen, White is the most generated race (47% of images), followed by Black (33%). The remaining races are rarer in comparison, e.g., 3% are Asian, and 5% are Indian. As for gender, males appear more frequently (65%). These findings mirror what was reported by Ghosh and Caliskan [11] who showed that most images generated by Stable Diffusion (v2.1) representing the prompt “a front-facing photo of a person” depict light-skinned Western men.
@@
One possible explanation behind these results could be that SDXL merely reflects the biases already present in the dataset on which it was trained, namely LAION-5B [27]. To examine this possibility, we used a subset of LAION-5B [28] consisting of 88,714 images; see Material and Methods for more details. The distribution of those images is summarized by the plain bars in Figure 1. As can be seen, images depicting White individuals are more frequent in LAION-5B than SDXL (63% vs. 47%), while those depicting other races are less frequent. In terms of gender, both men and women appear in LAION-5B with equal probability. This indicates that SDXL contains biases that cannot be fully explained by the data on which it was trained.
Figure 1: Examining biases in LAION-5B, Stable Diffusion XL (SDXL), and our SDXL-Inc. Comparing gender and race distributions in LAION-5B, SDXL, and our SDXL-Inc based on a sample of 88,714 images from the LAION-5B dataset, 10,000 images generated by SDXL, and 10,000 generated by SDXL-Inc. For the latter 20,000 images, we used the prompt: “a photo of a person”.
Some SDXL users may consider the racial distribution of images to be unsatisfactory, especially if it underrepresents certain groups compared to the society in which these users reside. One way to address this issue is to develop a solution that allows users to specify their desired distributions of race and gender. Perhaps the most intuitive target distribution is the one in which different groups are represented equally. With this in mind, we introduce such a debiasing solution, and test its ability to represent genders and races equally, although the same techniques can be used with any given target distribution. More specifically, we introduce a fine-tuned version of SDXL, which we call: “SDXL-Inc” (Inc stands for Inclusive).
@@
Our model was created as follows: First, we identified 32 professions and divided them into 21 for fine-tuning and 11 for testing (these professions are listed in Supplementary Table 1). Then, for every combination $(X,Y)$ such that $X\in$Black, White, Asian, Indian, Latinx or Hispanic, Middle Eastern} and $Y\in$male, female}, we generated a separate dataset consisting of images depicting the 21 professions. The generation of these 12 datasets (6 races × 2 genders) was carried out using the prompt: “a photo of $X$ $Y$ $Z$ looking at the camera, closeup headshot facing forward, ultra quality, sharp focus”, where $Z$ denotes one of the 21 professions. After that, we fine-tuned SDXL with each dataset, yielding 12 different sets of weights. The fine-tuning was done using LORA (Low Rank Adaptation) [29], which reduces the number of trainable parameters for the downstream task; see Materials and Methods for more details. The basic idea of SDXL-Inc is to randomly select one of those 12 sets of weights based on the target distribution of interest (which happens to be uniform in our experiments).
@@
Having observed that White males dominate SDXL-generated images, our next objective was to determine if this trend translates into specific professional stereotypes. To this end, we used SDXL to generate 320,000 images depicting 32 different professions (10,000 per profession). More specifically, for each profession, $Z$, we used the following prompt: “a photo of a $Z$, looking at the camera, closeup headshot facing forward, ultra quality, sharp focus”. Moreover, to specify the types of images that we want to avoid, we used the following negative prompt: “cartoon, anime, 3d, painting, b&w, low quality”. For each profession, we used our classifier to examine the racial- and gender- composition of the resulting images.
@@
We start off by examining racial stereotypes in each profession. Figure 2a depicts the results for the 25 professions, while Supplementary Figure 2 depicts the remaining seven professions. Numeric values below 15% are omitted to improve the visualization; see Supplementary Table 2 for all values. As can be seen, White is the most frequently generated race in 21 out of the 25 professions. As for the remaining four, two of them are among the least prestige occupations, namely Cleaner and Security Guard [30], and both are mostly represented by images depicting Black individuals. Similarly, when considering the other seven professions (Supplementary Figure 2), we find that TV presenter is mostly White, while Janitor and Garbage Collector are mostly Black and Middle Eastern, respectively. These findings confirm the findings of [9], showing that prestigious, high-paying professions are often represented as White.
Having examined the racial distribution of professions, we now shift our attention to the gender distribution. The results are depicted in Figure 2b; and the exact percentages of males and females are listed in Supplementary Table 2. As can be seen, males represent 90% of the images in 16 different professions, including Doctor and Professor—two highly prestigious occupations [30]. Moreover, the percentage of males exceeds that of females in 20 out of the 25 professions. As for the remaining five, they include Nurse and Secretary—two common stereotypes of women [31]. SDXL’s tendency to depict Nurses as predominantly female and Doctors as predominantly male echoes long-standing gender stereotypes in these professions [32]. Similar findings were reported in the literature, showing that AI-generated images tend to associate certain jobs with men and certain other jobs with women [13, 10].
The stereotypes documented in Figure 2 are alarming, as stereotypical portrayals have been shown to reinforce biases and limit ambitions [33, 34]. In contrast, portrayals of underrepresented groups in diverse professional roles can inspire similar career aspirations among these groups, e.g., seeing women in STEM roles can motivate young girls to pursue these fields [35].
Debiasing Stable Diffusion across professions and attributes
Building on the examination of professional stereotypes, this section expands the scope of analysis to encompass a broader set of attributes that are often subject to societal biases. In particular, we focus on the following attributes: Beautiful, Intelligent, Winner, Terrorist, Criminal, Poor, Parent, and Sibling. These attributes are deeply intertwined with social and cultural stereotypes. More specifically, “Beautiful” often carries implicit biases related to gender and race, with certain features being considered more desirable based on prevailing societal norms [36, 37]. Similarly, “Intelligent” and “Winner” are often associated with masculine traits [14], while attributes like “Terrorist”, “Criminal”, and “Poor” can trigger racial biases [38, 39]. Finally, “Parent” and “Sibling” were chosen as universal attributes that should, at least in principle, not be more strongly linked to one particular gender or race than another. Analyzing these attributes allows us to assess how SDXL portrays these complex social constructs.
The results of this analysis are depicted in the upper row of Figure 3a, and the corresponding numerical values are listed in Supplementary Table 5. We find that White dominates the three attributes that tend to be associated with success and attractiveness, i.e., Winner, Beautiful, and Intelligent. White also dominates the two family-related attributes, i.e., Parent and Sibling. In contrast, when it comes to Terrorism, Middle Eastern is the most common race, and none of the images depict a White individual, reinforcing existing stereotypes connecting Middle Easterners to terrorism [40]. Similarly, when it comes to crime and poverty, the majority of images depict Black individuals, reinforcing common stereotypes in this regard [41]. Similar findings were reported by Bianchi et al. [9] who analyzed Stable Diffusion (v1-4) and showed that attractiveness is associated with White people, thugs are associated with Black people, and terrorists are associated with Brown men with beards.
@@
Having established that SDXL exhibits biases in terms of attributes, we now evaluate SDXL-Inc’s ability to address these biases. To this end, we repeated the same procedure used earlier with SDXL, but using our SDXL-Inc model instead. The results for professions are depicted in the bottom row of Figure 3a. As shown, races are represented more uniformly compared to SDXL, as evidenced by the substantial reduction in standard deviation ($\sigma$). Importantly, none of the eight attributes considered in this analysis was used during the fine-tuning phase, and yet SDXL-Inc was able to significantly reduce the racial biases related to them. This indicates that our solution can indeed be generalized beyond the features it was fine-tuned on.
@@
Let us now reexamine professional stereotypes to further evaluate SDXL-Inc’s debiasing capability. To this end, we selected four of the professions that were used during the fine-tuning phase (Dietitian, Manager, Pharmacist, and Pilot), as well as four professions that were not used during that phase (Accountant, Journalist, Musician, and Firefighter) to further assess the generalizability of SDXL-Inc. For each profession, Figure 3b shows the racial distribution of images using SDXL (upper row) and SDXL-Inc (lower row). As can be seen, regardless of whether the profession is Black-dominated (Musician and Firefighter) or White-dominated, our solution is able to significantly reduce difference between races. Again, the reduction in standard deviation is substantial, reaching multiple folds across all professions. Similar trends are observed when examining the remaining 24 professions considered in our analysis; see Supplementary Figure 3a.
@@
d
[…]
Quantifying and addressing racial homogenization
Motivated by Edward Said’s concept of Orientalism, which critiques the West’s tendency to represent Eastern cultures and people through reductive and homogenous imagery [18], we examined the potential for SDXL to exhibit similar patterns of racial homogenization. Specifically, we investigated whether SDXL generated images of certain racial groups with overly similar facial features, echoing Orientalist representations that flatten the complexities of individual identities. To this end, we started off by manually examining a random sample of SDXL-generated images depicting Middle Eastern individuals. As shown in Figure 4a, such images depict men as being extremely similar to one another, e.g., they are all bearded and brown-skinned. Similarly, images representing women appear similar to each other, e.g., they have almost identical skin tones, and they all appear to be wearing traditional attire. While this random sample is indeed very small, it motivates a thorough examination of the phenomenon across racial and gender groups. Moreover, this sample highlights the need to develop a solution that increases the diversity of facial features when depicting individuals of a given race or gender.
To quantify the degree of racial homogenization in SDXL, we used it to generate 1,000 images per race, and obtained an embedding for each image using our racial classifier. Finally, for every image of a particular race, we measured the average cosine similarity between that image and all other images of that race. The result of this analysis is depicted in Figure 4b. As shown in this figure, the degree to which White individuals resemble one another is smaller than that of any given race. To put it differently, the facial diversity of White people is greater than that of other races. Interestingly, images depicting Middle Eastern individuals are the least diverse, which reinforces Orientalism.
@@
The above results motivate the development of a version of SDXL that generates images with greater facial diversity per race (compared to the original version). To this end, we downloaded from Flickr-Faces-HQ high-resolution (1024×1024) images of human faces with “considerable variation in terms of age, race, and image background” [43]. Since the dataset is unlabeled, we used our classifier to infer the race of each image; see Supplementary Table 4 for the number of images per race. The resulting labeled dataset was then used to fine-tune SDXL. Again, the fine-tuning was done using LORA [29] to reduce the number of trainable parameters for the downstream task. We call the resulting model “SDXL-Div” (where Div is a shorthand for Diversity); see Materials and Methods for technical details. To evaluate this model, we used it to generate 1,000 images and for each one of them, we measured the average cosine similarity to all other images of that race.
@@
As can be seen in Figure 4c, SDXL-Div is indeed able to increase the facial diversity of SDXL, regardless of the race. The greatest difference is observed in the case of Middle Eastern individuals (dropping the mean cosine similarity from 0.61 to 0.41) and Latinx individuals (from 0.54 to 0.39). To help the reader appreciate the difference in diversity between the two models, Figure 4d provides sample images of Middle Eastern men and women generated SDXL-Div. As can be seen, images generated using this model have markedly greater facial diversity compared to those generated using the former model.
Survey experiments
@@
Given the increasing pervasiveness of AI-generated images, it is crucial to understand how repeated exposure to such images may shape public perceptions, especially concerning sensitive topics like race and gender. To this end, we ran four survey experiments. We estimated that to obtain a power of 0.8 to detect a medium effect size (Cohen’s $d$) of 0.5 in a paired-sample comparison, a sample of 135 participants would be needed. We recruited participants on Prolific, with prescreening settings of US residents and English as native language. Additionally, we prohibited participation more than once in our experiment. These survey experiments were preregistered at AsPredicted (https://aspredicted.org/3v3wt.pdf). All experiments were approved by the Institutional Review Board at New York University Abu Dhabi under the category of Exempt or Expedited Research (HRPP-2024-53).
@@
Study 2 examines gender bias, and is similar to Study 1 apart from two modifications: (i) The professions are {accountant,math $\{\textnormal{accountant},\textnormal{math scientist},\textnormal{and}\textnormal{tailor}\}$; (ii) after seeing all six images, participants answer the question $Q_{2}$: What percentage of $P$ in the US are men? Here, the non-inclusive images are SDXL-generated, and all depict men, while the inclusive images are generated by SDXL-Inc and half depict men while the other half depicts women. In both conditions, each of the six races is depicted in a separate image; see Supplementary Figure 9.